OpenStudy (anonymous):

In order to sum a geometric sequence with seven terms, how would the equation be set up?

OpenStudy (bibby):

do you have the equation to sum up X amount of terms in a geometric series?

OpenStudy (anonymous):

No I do not, that is what I was inquiring about.

OpenStudy (bibby):

\[\huge S_n=a_1(\frac{{1-(r}^n)}{1-r})\]

OpenStudy (anonymous):

Thank you bibby!

OpenStudy (bibby):

so n is the amount of terms, can you take it from here?

OpenStudy (anonymous):

Correct. My series is -4, 24, -144 I know the common ratio is -6. So i am plugging in the values now.

OpenStudy (anonymous):

It is a 7 term series

OpenStudy (anonymous):

Actually, isnt the equation set up like this?

OpenStudy (bibby):

I don't think so

OpenStudy (bibby):

If you factor out the \(a_1\) from the numerator, it's the same thing

OpenStudy (anonymous):

Interesting because when I set up the equation I get a decimal. And not a whole number.

OpenStudy (bibby):

I'll step through it then

OpenStudy (bibby):

\[\huge S_n=a_1(\frac{{1-(r}^n)}{1-r})\] \(\huge S_n=-4(\frac{{1-(-6}^7)}{1-(-6)})=-4(\frac{{1--279936})}{1+6})\)

OpenStudy (anonymous):

Oh wow. Nevermind. I messed up. I left out the "r" on the bottom so i was ending up with -1119748. Instead of -159964. Thank you very much! And sorry to take up lots of your time :/

OpenStudy (bibby):

lol, I have a bazillion sites I would have been visiting anyways, I was tabbed out, you were doing most of the work

OpenStudy (anonymous):

Still, Thank you!

OpenStudy (bibby):

np np :)